ISRAEL JOURNAL OF MATHEMATICS

TABLE OF CONTENTS, VOLUME 249, 2022

André, S. Homogeneity in virtually free groups  167

Ardizzoni, A., P. Saracco and D. Ştefan PBW-deformations of graded rings

769

Ballmann, W. and P. Polymerakis Equivariant discretizations of diffusions and harmonic functions of bounded growth

227

Banakh, T., R. Bonnet and W. Kubiś Vietoris hyperspaces over scattered Priestley spaces

37

Behajaina, A., B. Deschamps and F. Legrand Problèmes de plongement finis sur les corps non commutatifs

617

Bell, J. P. and P. V. Danchev Affine representability and decision procedures for commutativity theorems for rings and algebras

121

Biswas, A. and J. P. Saha Spectral bounds of directed Cayley graphs of finite groups

973

Bonnet, R. See T. Banakh Cai, F. and S. Shao Topological characteristic factors and indepen-

dence along arithmetic progressions

259

Danchev, P. V. See J. P. Bell Deng, H., H. Liu and X. Yang Critical points of solutions to a kind

of linear elliptic equations in multiply connected domains

935

Deschamps, B. See A. Behajaina

Friedman, S.-D. and M. Levine Patterns of stationary reflection  1
Hawthorne, C. Automata and tame expansions of (ℤ, +)  651

Kim, M. and A. Lew Complexes of graphs with bounded independence number

83

Klin, M. See M. Muzychuk Kubiś, W. See T. Banakh Legrand, F. See A. Behajaina Levine, M. See S.-D. Friedman Lew, A. See M. Kim Lim, M. F. On the growth of even K-groups of rings of integers in

p-adic Lie extensions

735

Liu, H. See H. Deng

Măntoiu, M. On Persson’s formula: an étale groupoid approach  899

Montaldo, S., C. Oniciuc and A. Ratto Polyharmonic hypersurfaces into space forms

343

Moretó, A. and G. Navarro Characters, exponents and defects in p-solvable groups

553

Moschetti, R. and G. P. Pirola Hyperelliptic odd coverings  477

Müller, S., P. Schlicht, D. Schrittesser and T. Weinert Lebesgue’s density theorem and definable selectors for ideals

501

Muzychuk, M., S. Reichard and M. Klin Jordan schemes  309

Navarro, G. See A. Moretó Nguyen, D.-M. Topological Veech dichotomy and tessellations of the

hyperbolic plane

577

Oniciuc, C. See S. Montaldo Orr, M., A. N. Skorobogatov, D. Valloni and Yu.G. Zarhin

Invariant Brauer group of an abelian variety

695

Pirola, G. P. See R. Moschetti Polymerakis, P. See W. Ballmann Porat, G. Lubin–Tate theory and overconvergent Hilbert modular

forms of low weight

431

Ratto, A. See S. Montaldo Reichard, S. See M. Muzychuk Saha, J. P. See A. Biswas Saracco, P. See A. Ardizzoni Schlicht, P. See S. Müller Schrittesser, D. See S. Müller Shao, S. See F. Cai Shi, Y., F. Yang and J. Yang A countable partition for singular

flows, and its application on the entropy theory

375

Skorobogatov, A. N. See M. Orr Spinka, Y. Finitary isomorphisms of renewal point processes and

continuous-time regenerative processes

857

Ştefan, D. See A. Ardizzoni Valloni, D. See M. Orr Weinert, T. See S. Müller Yang, F. See Y. Shi Yang, J. See Y. Shi Yang, X. See H. Deng Zarhin, Yu. G. See M. Orr